# NL2KQL: From Natural Language to Kusto Query

### v1

@@ Abstract.
Data is growing in volume and complexity, in turn, becoming harder to navigate even for the query language experts. Query languages are featureful and powerful; however they present a learning curve that may not be readily mastered by all users. Reducing technical obstacles would enable more data enthusiasts to access, understand and benefit from the data; hence, there is an ever increasing research on Natural Language (NL) Interface for Databases (NLIDB) to empower both technical and non-technical users to pose queries in NL and communicate with data. With Structured Query Language (SQL) being the most widely used query language, many NLQ-to-SQL solutions have been proposed over the years ranging from parsing-based approaches, to more recent Neural Machine Translation (NMT) techniques (Katsogiannis-Meimarakis and Koutrika, 2023). The two widely used NLQ-SQL datasets, WikiSQL (Zhong et al., 2017) and Spider (Yu et al., 2018b), have led to nearly a thousand NLQ-to-SQL solutions.
@@
In this study, we introduce an innovative approach for transforming Natural Language Queries (NLQ) into Kusto Query Language (KQL) queries, denoted as NL2KQL. This solution capitalizes on the capabilities of LLMs. The generated KQLs are semantically valid for the target Kusto cluster using cluster-specific synthetic few-shots and a Semantic Data Catalog. The principal contributions of this research are multifold. Firstly, we present the first end-to-end automated NLQ-to-KQL framework, incorporating novel methods for cluster-specific schema refinement and on-the-fly query refinement. Secondly, we introduce a method to synthesize large scale NLQ-KQL pairs with high variance tuned for a Kusto cluster. Thirdly, we propose a set of offline and online metrics to assess the effectiveness of our approach, some of which hold the potential for application in areas beyond KQL generation. Fourthly, we release the first benchmark for KQL generation evaluation, along with their Semantic Data Catalog, comprising 400 NLQ-KQL pairs compiled by KQL experts, distributed across two Kusto clusters and segmented into easy and hard categories. Lastly, an exhaustive ablation study is conducted to delineate each component’s individual contributions to the overall performance.
@@ 2.1. Kusto Query Language (KQL)
KQL, albeit simple, is a powerful data analytics tool with a wide range of aggregation functions  [^2] https://learn.microsoft.com/en-us/azure/data-explorer/kusto/query/aggregation-functions. The syntax is flexible and support various other syntactical forms and constructs that do not strictly adhere to the above mentioned pipeline pattern. These query forms are further discussed in Appendix A.
@@ 2.2. NLQ-to-KQL Problem Statement
The NLQ-to-KQL problem is formalized as: Given a Natural Language Query (NLQ) and a Kusto cluster ($c$) with a known schema ($s$), generate a KQL query ($\hat{q}$) which is valid with respect to $s$ and, if executed, fetches results that match the user’s intent.
@@ 2.3. NLQ-to-KQL Challenges
On the other hand KQL is not forgiving when it comes to syntactic or semantic errors. The generative model has to have knowledge of the Kusto cluster schema which may not be well-structured. KQL is designed for logs, time-series and telemetry, focusing on data exploration and analytics. KQL supports joining tables based on matching values, however, unlike SQL, it has no notion of primary and foreign keys and does not enforce referential integrity or uniqueness constraints. Complicating matters, commonly, KQL queries parse JSON columns to extract keys and the extracted JSON key, in turn, can be used to join tables (see Appendix A.7). Furthermore, telemetry data is unnormalized, inherently redundant, with columns duplicated across tables, all of which are confusing to an auto-regressive generative language model.
@@ 3.1. Semantic Data Catalog
A Semantic Data Catalog plays a pivotal role in encapsulating information about the structure, semantics, and contextual attributes of a database. The Schema Refiner and Query Refiner both leverage Data Catalog and the embeddings of its elements to make decisions for schema filtering and query repair.
@@
The Semantic Data Catalog capture both structural and semantic information and is created via an automated process. For a given Kusto cluster, the schema structure, including table names, column names, and column types are retrieved, either directly from the cluster (.show database schema) or a third-party API. Once the structure is retrieved, available public documents are extracted from Azure Monitor Data Reference  [^3] https://learn.microsoft.com/en-us/azure/azure-monitor/reference/ including table descriptions, column descriptions, and in some cases, value descriptions (e.g. description of error codes). Generic data retrieval queries are executed on the Kusto cluster to obtain a sample of data for each table. This data is used as a proxy to extract unique values for the low-cardinality Enum columns. The acuiqred descriptions ensure that the Semantic Data Catalog contains both the structure and the semantics of the elements in database.
@@
The embeddings of the extended summaries of all elements are calculated via a similarity embedding model and stored in a vector database for use at inference time. The extended summary is particularly helpful when dealing with columns like ”ErrorCode” which contain values numerical values (e.g. 16000 and 40008) necessitating contextual information for accurate understanding.
@@ 3.2. Schema Refiner
In response to these challenges, the Schema Refiner selects most relevant tables, columns, and potential values, and incorporates them in model’s context. The refiner relies on similarity of NLQ embeddings with elements of the Data Catalog.
@@ 3.3. Synthetic Few-shot Database
The NL2KQL solution uses a synthetic few-shot database (FSDB) to provide guidance for the LLM. The FSDB is synthesized per Kusto cluster and contains KQLs which are valid with respect to the cluster’s schema. The generated KQLs go through 3 rounds of validation to ensure sanity, syntactic and semantic correctness, and to filter out any ambiguous NLQs.
As depicted in Figure 2, the process starts with the Kusto cluster schema. A random table is sampled from the schema. For 30% of queries, another random table is chosen to join with the first table. When choosing the joint table, in 50% of occurrences the table is only selected from the subset of tables which have similar column types. The final set of tables (i.e., one table without join, and two tables with join) are passed to the next step.
In the second phase, a theme is randomly sampled from the following five themes: 1) Explore: Look for signs or hints of a security attack; 2) Expansion: Searches for additional contextual understanding for the scenario; 3) Detect: Look for events related to a security attack; 4) Remediate: Identify all evens for a given entity or asset; 5) Report: Provide summary statistics that helps with writing a report. The inclusion of random themeses further promotes the variability of the final KQLs.
In the third phase, an LLM is prompted to generate a KQL query given the table set and the theme. The prompt contains instructions on Kusto syntax with query design options. The prompt also makes suggestions regarding the difficulty of the query and how to construct the final KQL including number of pipes.
@@ 3.4. Few-shot Selector
Considering the above, a few-shot database (FSDB) is created where each sample is an NLQ-KQL pair. The embeddings of the NLQ part of all few-shots are calculated offline and stored in a vector database for use at inference time. During inference, given a user NLQ, the FSDB is primarily filtered to only include schema-relevant few-shots. The embeddings of the filtered few-shots are then compared with the embeddings of user’s NLQ and the top $f$ few-shots are selected based on their cosine similarity.
@@ 4. EXPERIMENTS
Moreover, to create a more challenging environment with a reasonably large schema size, it was assumed that the user has access to all the tables in the Kusto cluster’s schema. During Schema Refinement, maximum number of tables, $t$, was set to 9. Maximum number of values per each Enum column, $v_{\text{n}}$, was set 5. The number of few-shots, $f$, was set to 5.
@@ 4.1. Evaluation Benchmark
The proposed NL2KQL solution was evaluated on two Kusto clusters with different schemas.
Sentinel: The first schema originated from a Kusto cluster of Microsoft Sentinel, a Security Information and Event Management (SIEM) solution that delivers intelligent security analytics and threat intelligence across the enterprise. The Sentinel cluster has 23 tables. A team of security operations center analysts drafted 189 NLQ-KQL pairs targeted at the Sentinel cluster which are entirely used as the evaluation set.
Defender: The second schema is from Microsoft Defender XDR, an Extended Detection and Response (XDR) solution that coordinates detection, prevention, investigation, and response across endpoints, identities, email, and applications pre- and post-breach. The Defender cluster has 29 tables. Following the same process as the other cluster, 224 NLQ-KQL pairs were drafted as the Defender evaluation set.
In both clusters, the security analysts were tasked to write NLQ-KQL pairs that represent their day-to-day queries without including any Personally identifiable information (PII). The analysts had zero exposure to the Few-Shot Databases to mitigate any unintended bias and information leakage.
The two evaluation sets for Defender and Sentinel, along with their respective Semantic Data Catalogs, are published and publicly accessible at https://github.com/micro¡to-be-released-post-acceptance¿. A summary of the schema, including list of tables and columns are available in Appendix C.
@@ 4.2.2. Online Metrics
The inherent data redundancy of logs and telemetries in Kusto clusters and KQL’s flexibility in expressing the same query in different ways erodes uniqueness of ground-truth KQL that corresponds to a given NLQ. Online metrics, which measure the similarities of the query results, are intended to address the limitations of offline metrics, which only evaluate the syntactic and semantic similarity of the queries. The online metrics are as followed.
@@ 5. Results and Analysis
Results show that removing Query Refiner has a negative impact on both syntactic and semantic scores, reducing them to $0.9875$ and $0.9449$, respectively, which, in turn, negatively impacts the $\text{Avg}_{\text{exec}}$ metric as queries with errors fail to execute. Despite its positive impact, the Query Refiner made a mistake on two samples in the evaluation set slightly impacting the $\text{Filter}_{\text{col}}$ value of the NL2KQL method. Query Refiner is a deterministic post-processing component, therefore, we closely studied its behaviour on the evaluation set and observed that, on average it fixed 8.6 queries in the Defender cluster and 4.3 queries in the Sentinel cluster across the triplicate runs. Details of this study is reported in Appendix B.
@@ 6. Discussion And Future Work
Among future works, we envision the trending integration of LLMs as agent-like entities for iterative interaction and error resolution to be a viable option. In an agent-like setup, an initially generated KQL query would engage the cluster to fetch some samples. This process will leverage both the results and any exceptions encountered as a basis for iteration and debugging. A significant challenge associated with this agent-based iterative methodology is the increase of response times, shifting interactions away from real-time and adversely affecting the user experience. Similarly, the Query Refiner component could have been implemented to involve instructing an LLM to regenerate the query based on the diagnostic code from the Kusto parser. While effective, this approach would have introduces additional latency due to the slow auto-regressive nature of decoder-only LLMs. Given the real-time nature of NL2KQL, response time is a concern in design decisions.
Finally, extending the NL2KQL to a chat-based text interface enables users to inquire further about the data through follow-up questions. This process generates a new KQL that incorporates the augmented context. Although this approach presents new challenges, it significantly enhances the accessibility of Kusto clusters for data analysts, facilitating more intuitive interactions.
@@ Acknowledgements.
Appendix A KQL Forms
@@ A.2. Subqueries
A.3. Scalar and Tabular Functions
@@ A.6. Conditional Statements
A.7. JSON Parse and Join
@@
Appendix B Query Repair Examples
@@
Appendix C Cluster Schema
C.1. Defender Cluster

### v2

@@ 1. Introduction
Data is growing in volume and complexity, in turn, becoming harder to navigate even for the query language experts. Query languages are featureful and powerful; however they present a learning curve that may not be readily mastered by all users. Reducing technical obstacles would enable more data enthusiasts to access, understand and benefit from the data; hence, there is an ever increasing research on Natural Language (NL) Interface for Databases (NLIDB) to empower both technical and non-technical users to pose queries in NL and communicate with data. With Structured Query Language (SQL) being the most widely used query language, many NLQ-to-SQL solutions have been proposed over the years ranging from parsing-based approaches, to more recent Neural Machine Translation (NMT) techniques (Katsogiannis-Meimarakis and Koutrika, 2023). The two widely used NLQ-SQL datasets, WikiSQL (Zhong et al., 2017) and Spider (Yu et al., 2018b), and the recently released BIRD benchmark (Li et al., 2023), have led to nearly a thousand NLQ-to-SQL solutions.
@@
In this study, we introduce an innovative approach for transforming Natural Language Queries (NLQ) into Kusto Query Language (KQL) queries, denoted as NL2KQL. This solution capitalizes on the capabilities of LLMs. The generated KQLs are semantically valid for the target Kusto database using database-specific synthetic few-shots and a Semantic Data Catalog. The principal contributions of this research are multifold. Firstly, we present the first end-to-end automated NLQ-to-KQL framework, incorporating novel methods for database-specific schema refinement and on-the-fly query refinement. Secondly, we introduce a method to synthesize large scale NLQ-KQL pairs with high variance tuned for a Kusto database. Thirdly, we propose a set of offline and online metrics to assess the effectiveness of our approach, some of which hold the potential for application in areas beyond KQL generation. Fourthly, we release the first benchmark for KQL generation evaluation, along with their Semantic Data Catalog, comprising 400 NLQ-KQL pairs compiled by KQL experts, distributed across two Kusto databases and segmented into easy and hard categories. Lastly, an exhaustive ablation study is conducted to delineate each component’s individual contributions to the overall performance.
@@ 3.1. Kusto Query Language (KQL)
Similar to SQL databases, a kusto database is defined by its schema which contains the names of tables and their respective column names and column types.
@@
KQL, albeit simple, is a powerful data analytics tool with a wide range of aggregation functions  [^2] https://learn.microsoft.com/en-us/azure/data-explorer/kusto/query/aggregation-functions. The syntax is flexible and support various other syntactical forms and constructs that do not strictly adhere to the above mentioned pipeline pattern. These query forms are further discussed in Appendix B.
@@ 3.2. NLQ-to-KQL Problem Statement
The NLQ-to-KQL problem is formalized as: Given a Natural Language Query (NLQ) and a Kusto database with a known schema ($s$), generate a KQL query ($\hat{q}$) which is valid with respect to $s$ and, if executed, fetches results that match the user’s intent.
@@ 3.3. NLQ-to-KQL Challenges
On the other hand KQL is not forgiving when it comes to syntactic or semantic errors. The generative model has to have knowledge of the Kusto database schema which may not be well-structured. KQL is designed for logs, time-series and telemetry, focusing on data exploration and analytics. KQL supports joining tables based on matching values, however, unlike SQL, it has no notion of primary and foreign keys and does not enforce referential integrity or uniqueness constraints. Complicating matters, commonly, KQL queries parse JSON columns to extract keys and the extracted JSON key, in turn, can be used to join tables (see Appendix B.7). Furthermore, telemetry data is unnormalized, inherently redundant, with columns duplicated across tables, all of which are confusing to an auto-regressive generative language model.
@@ 4.1. Semantic Data Catalog
A Semantic Data Catalog plays a pivotal role in encapsulating information about the structure, semantics, and contextual attributes of a database. The Schema Refiner and Query Refiner both leverage Data Catalog and the embeddings of its elements to make decisions for schema filtering and query repair (Figure 3).
@@
The Semantic Data Catalog capture both structural and semantic information and is created via an automated process. For a given Kusto database, the schema structure, including table names, column names, and column types are retrieved, either directly from the database (.show database schema) or a third-party API. Once the structure is retrieved, available public documents are extracted from Azure Monitor Data Reference  [^3] https://learn.microsoft.com/en-us/azure/azure-monitor/reference/ including table descriptions, column descriptions, and in some cases, value descriptions (e.g. description of error codes). Generic data retrieval queries are executed on the Kusto database to obtain a sample of data for each table. This data is used as a proxy to extract unique values for the low-cardinality Enum columns. The acquired descriptions ensure that the Semantic Data Catalog contains both the structure and the semantics of the elements in database.
@@
The embeddings of the extended summaries of all elements are calculated via a similarity embedding model (Figure 3) and stored in a vector database for use at inference time. The extended summary is particularly helpful when dealing with columns like ”ErrorCode” which contain values numerical values (e.g. 16000 and 40008) necessitating contextual information for accurate understanding.
@@ 4.2. Schema Refiner
In response to these challenges, and similar to the Schema Linking practices of NL-to-SQL (Zhang et al., 2024), the Schema Refiner selects most relevant tables, columns, and potential values, and incorporates them in model’s context. The refiner relies on similarity of NLQ embeddings with elements of the Data Catalog.
@@ 4.3. Synthetic Few-shot Database
The NL2KQL solution uses a synthetic few-shot database (FSDB) to provide guidance for the LLM. The FSDB is synthesized for a target Kusto database and contains KQLs which are valid with respect to the database’s schema. The generated KQLs go through 3 rounds of validation to ensure sanity, syntactic and semantic correctness, and to filter out any ambiguous NLQs.
The synthetic KQL generation starts with random table sampling from the database schema. For 70% of queries only one table is sampled. For the remaining 30%, another random table is chosen to join with the first table. When choosing the joint table, in 50% of occurrences the table is only selected from the subset of tables which have similar column types. The final set of tables (i.e., one table without join, and two tables with join) are passed to the next step (Figure 2).
Similarly, a theme is randomly sampled from the following five themes: 1) Explore: Look for signs or hints of a security attack; 2) Expansion: Searches for additional contextual understanding for the scenario; 3) Detect: Look for events related to a security attack; 4) Remediate: Identify all evens for a given entity or asset; 5) Report: Provide summary statistics that helps with writing a report. The inclusion of random themes further promotes the variability of the final KQLs.
An LLM is then prompted to generate a KQL query given the table set (schema) and the theme (Figure 2). The prompt contains instructions on Kusto syntax with query design options. The prompt also makes suggestions regarding the difficulty of the query and how to construct the final KQL including number of pipes.
@@ 4.4. Few-shot Selector
Considering the above, a few-shot database (FSDB) is created where each sample is an NLQ-KQL pair. The embeddings of the NLQ part of all few-shots are calculated offline and stored in a vector database for use at inference time (Figure 3). During inference, given a user NLQ, the FSDB is primarily filtered to only include schema-relevant few-shots. The embeddings of the filtered few-shots are then compared with the embeddings of user’s NLQ and the top $f$ few-shots are selected based on their cosine similarity.
@@ 5. EXPERIMENTS
Moreover, to create a more challenging environment with a reasonably large schema size, it was assumed that the user has access to all the tables in the Kusto database’s schema. During Schema Refinement, maximum number of tables, $t$, was set to 9. Maximum number of values per each Enum column, $v_{\text{n}}$, was set 5. The number of few-shots, $f$, was set to 5.
@@ 5.1. Evaluation Benchmark
The proposed NL2KQL solution was evaluated on two Kusto databases with different schemas.
Sentinel: The first schema originated from a Kusto database of Microsoft Sentinel, a Security Information and Event Management (SIEM) solution that delivers intelligent security analytics and threat intelligence across the enterprise. The Sentinel database has 23 tables. A team of security operations center analysts drafted 189 NLQ-KQL pairs targeted at the Sentinel database which are entirely used as the evaluation set.
Defender: The second schema is from Microsoft Defender XDR, an Extended Detection and Response (XDR) solution that coordinates detection, prevention, investigation, and response across endpoints, identities, email, and applications pre- and post-breach. The Defender database has 29 tables. Following the same process as the other database, 224 NLQ-KQL pairs were drafted as the Defender evaluation set.
In both databases, the security analysts were tasked to write NLQ-KQL pairs that represent their day-to-day queries without including any Personally identifiable information (PII). The analysts had zero exposure to the Few-Shot Databases to mitigate any unintended bias and information leakage.
A summary of the schema, including list of tables and columns are available in Appendix D.
@@ 5.2.2. Online Metrics
The inherent data redundancy of logs and telemetries in Kusto databases and KQL’s flexibility in expressing the same query in different ways erodes uniqueness of ground-truth KQL that corresponds to a given NLQ. Online metrics, which measure the similarities of the query results, are intended to address the limitations of offline metrics, which only evaluate the syntactic and semantic similarity of the queries. The online metrics are as followed.
@@ 6. Results and Analysis
Results show that removing Query Refiner has a negative impact on both syntactic and semantic scores, reducing them to $0.9875$ and $0.9449$, respectively, which, in turn, negatively impacts the $\text{Avg}_{\text{exec}}$ metric as queries with errors fail to execute. Despite its positive impact, the Query Refiner made a mistake on two samples in the evaluation set slightly impacting the $\text{Filter}_{\text{col}}$ value of the NL2KQL method. Query Refiner is a deterministic post-processing component, therefore, we closely studied its behaviour on the evaluation set and observed that, on average it fixed 8.6 queries in the Defender database and 4.3 queries in the Sentinel database across the triplicate runs. Details of this study is reported in Appendix C.
@@ 7. Discussion And Future Work
Among future works, we envision the trending integration of LLMs as agent-like entities for iterative interaction, refinement and error resolution to be a viable option. In an agent-like setup samples are iteratively fetched from candidate tables to extend the model’s context and encountered exceptions are used as a basis for debugging. A significant challenge associated with this agent-based iterative methodology is the increase of response times, shifting interactions away from real-time. Similarly, the Query Refiner component could have been implemented to involve instructing an LLM to regenerate the query based on the diagnostic code from the Kusto parser. While effective, this approach would have introduces additional latency due to the slow auto-regressive nature of decoder-only LLMs. Given the real-time nature of NL2KQL, response time is a concern in design decisions.
Finally, extending the NL2KQL to a chat-based text interface enables users to inquire further about the data through follow-up questions. This process generates a new KQL that incorporates the augmented context. Although this approach presents new challenges, it significantly enhances the accessibility of Kusto databases for data analysts, facilitating more intuitive interactions.
@@ Acknowledgements.
Appendix A Prompt Template
Each section of the prompt fill its respective {{NAME_PLACEHOLDER}} placeholder in the following prompt template:
# Instructions
You will receive a natural language request. As a cyber-security expert specializing in Microsoft USX, your task is to write a Kusto Query Language (KQL) query that fulfills the request while adhering to all best practices.
Follow these steps:
- Start by coming up with a stepwise plan on how to address the request.
- Identify any challenges you might face with this plan.
- Choose a final approach that seems to fit the request best.
- Write an optimized and performant Kusto query that fulfills the request while adhering to all best practices.
# Schema
To help you write your query, you have been given the following Microsoft USX database schema.
Use your own discretion to decide which tables and columns are needed.
{{SCHEMA_PLACEHOLDER}}
# Possible Values
In addition, the user has identified the following values which can appear in these tables.
Use your own discretion whether to include any of these values and how best to use them.
Do not use any values that are not explicitly mentioned in the prompt or the section below.
{{VALUES_PLACEHOLDER}}
# Kusto Syntax
A sample of useful KQL syntax are listed below.
**Scalar Functions**
| Syntax | Usage |
| around(value, center, delta)‘ | Check if value is in the range center +/- delta |
| bin(value, roundto)‘ | Round value to the nearest multiple of roundto |
| isempty(value)‘ | Check if a column is null or empty |
| parse_json(string)‘ | Convert a JSON-like string into a dynamic object |
| todouble(value)‘ | Convert a value to a double |
| toint(value)‘ | Convert a value to an integer |
| toscalar(expression)‘ | Convert a single-row, single-column tabular expression into a scalar value |
| tostring(value)‘ | Convert a value to a string |
| range ( start, stop[, step ])‘ | Create a dynamic array of equally spaced values |
**Aggregation Functions**
| Syntax | Usage |
| arg_min(minExpr, returnExpr [, …])‘ | Find the row that minimizes an expression |
| count()‘ | Count the number of records in a table or group |
| count_distinct(expr)‘ | Count the number of unique values of expr per group |
| dcount(expr)‘ | Estimate the number of unique values of expr per group |
| make_set(expr)‘ | Create an array of unique values of expr per group |
| percentile(expr, percentile)‘ | Estimate the nearest-rank percentile of the population defined by expr |
**Window Functions**
| Syntax | Usage |
| row_number()‘ | Get the index of each row |
| next(column[, offset, default_value ])‘ | Get the value of an upcoming row |
| prev(column[, offset, default_value ])‘ | Get the value of a previous row |
**Tabular Operators**
| Syntax | Usage |
| distinct ColumnName[, ColumnName2, …]‘ | Produce a table of unique combinations of the input columns |
| mv-apply Columns [to typeof(DataType)] on ( SubQuery )‘ | Apply a subquery to each record and union the results |
| mv-expand [bagexpansion=(bag | array)] [Name =] Expr [to typeof(DataType)]‘ | Explode an array or JSON object into multiple rows |
| render Visualization | Render a visualization |
| serialize | Enable window functions |
**Scalar Operators**
| Syntax | Usage |
| contains ( string )‘ | Check if a value contains a substring |
| has ( string )‘ | Check if a value contains a specific word or term |
| has_any ( string, )‘ | Check if a value contains any substring in a set |
| has_all ( string, )‘ | Check if a value contains all substrings in a set |
| in~ ( string, )‘ | Check if a value equals any substring in a set |
| in ( number, )‘ | Check if a value equals any number in a set |
| matches regex string | Check if a value matches a regex pattern |
# Examples
{{EXAMPLES_PLACEHOLDER}}
# Reminder
All steps must be completed in a single message.
Remember, your output will contain queries like
~~~kusto
KQL QUERY GOES HERE
{{USER_REQUEST_PLACEHOLDER}}
Appendix B KQL Forms
@@ B.2. Subqueries
B.3. Scalar and Tabular Functions
@@ B.6. Conditional Statements
B.7. JSON Parse and Join
@@
Appendix C Query Repair Examples
@@
Appendix D Database Schema
D.1. Defender Database